OpenStudy (anonymous):

evaluate the given integral 3(e^x/4 + e^-x/4)dx

OpenStudy (anonymous):

is it this integral??? \(\large \int3(e^{x/4} + e^{-x/4})dx \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u can break it into 2 terms and intergrate them seperately...

OpenStudy (anonymous):

so it should look like \[3*e^{x/4} + 3*e^{-x/4}\]

OpenStudy (anonymous):

do u howto integrate something with and e in it?

OpenStudy (anonymous):

Are you sure it isn't \[ \int 3\left(\frac{e^x}{4}+\frac{e^{-x}}{4}\right)\;dx \]

OpenStudy (anonymous):

If it is indeed \[ \int 3(e^{x/4}+e^{-x/4})\;dx \]Then first do a \(u=x/4\) sub.